angle a and angle b form a linear pair. the measure of angle b is 10 less than twice the measure of angle a. find the measure of each angle
Found 2 solutions by nerdybill, themathprof:
Answer by nerdybill(7384)

angle a and angle b form a linear pair. the measure of angle b is 10 less than twice the measure of angle a. find the measure of each angle
.
Let a = measure of angle a
and b = measure of angle b
.
Because the form a "linear pair" the sum of the two angles must be 180:
a + b = 180 (equation 1)
.
From: "the measure of angle b is 10 less than twice the measure of angle a"
b =2a-10
-2a + b = -10 (equation 2)
.
Our equations:
a + b = 180
-2a + b = -10
.
Using elimination method:
multiply equation 2 by a -1:
a + b = 180
2a - b = 10
-------------
3a = 170
a = 56.667 deg (angle a)
or 56 2/3 deg
.
Substitute this back into equation 2 and solve for b:
a + b = 180
56.667 + b = 180
b = 180 - 56.667
b = 123.333 deg (angle b)
or 123 1/3 deg

Answer by themathprof(21)

angle a and angle b form a linear pair. the measure of angle b is 10 less than twice the measure of angle a. find the measure of each angle
I guess that you mean they form a straight angle which = 180 (degrees)
Look for the word 'than'. What's after the word 'than?' Angle a.
So, x=angle A
Now, B is 10 less
So B=x-10
But a straight angle has 180 degrees
So A+B=180
Or x+x-10=180
or, 2x-10=180
2x=190
x=95 (angle A)
x-10=(95-10)=85 (Angle B)
Find more word problems worked out
step by step in my FREE software
math911. Download and activate FREE
from www.math911.com
Tell your friends (and your teacher)
Professor Martin Weissman